# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 2.5. Hopf $\infty$-operads
$\Gamma:\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})_{/{\mathbb{1}}}^{\prime}\to\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nu}}}({\mathcal{C}})$
that takes the fiber of the augmentation in $\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nu}}}({\mathcal{C}}).$ For any ${\mathrm{X}}\in\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nu}}}({\mathcal{C}})$ the unit
@@
where the vertical functors are fully faithful and the bottom functor is an equivalence. Since ${\mathcal{C}}\subset{\mathcal{D}}$ is closed under retracts, ${\mathrm{A}}\in\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nu}}}({\mathcal{D}})$ belongs to $\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nu}}}({\mathcal{C}})$ if its image $\bar{{\mathrm{F}}}({\mathrm{A}})\simeq{\mathrm{A}}\oplus{\mathbb{1}}$ belongs to $\mathrm{Alg}_{\mathrm{Comm}}({\mathcal{C}})_{/{\mathbb{1}}}$. So the top functor is an equivalence.
@@
    There is a canonical embedding ${\mathrm{coAlg}}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\hookrightarrow{\mathrm{coAlg}}_{{\mathrm{coComm}}}({\mathcal{C}})_{{\mathbb{1}}/}$ sending ${\mathrm{X}}$ to ${\mathrm{X}}\oplus{\mathbb{1}}.$
@@
    If ${\mathcal{C}}$ admits cofibers, the embedding ${\mathrm{coAlg}}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\hookrightarrow{\mathrm{coAlg}}_{{\mathrm{coComm}}}({\mathcal{C}})_{{\mathbb{1}}/}$ admits a left adjoint that takes the cofiber of the coaugmentation.
@@
    The embedding ${\mathrm{coAlg}}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\hookrightarrow{\mathrm{coAlg}}_{{\mathrm{coComm}}}({\mathcal{C}})_{{\mathbb{1}}/}$ is an equivalence if ${\mathcal{C}}$ is stable or idempotent complete and additive.
The next proposition is [20, Theorem 8.9.]:
@@
We obtain the following relationship between $\mathrm{Comm}$-algebras and ${\mathbb{E}}_{\infty}$-algebras:
@@
$\mathrm{Sym}:=\mathrm{Comm}\circ(-):{\mathcal{C}}\to\mathrm{Alg}_{\mathrm{Comm}}({\mathcal{C}}),\ {\mathrm{X}}\mapsto\mathrm{Comm}\circ{\mathrm{X}}\simeq\coprod_{{\mathrm{n}}\geq 0}({\mathrm{X}}^{\otimes{\mathrm{n}}})_{\Sigma{\mathrm{n}}}.$
[…]
@@
[…]
@@
By [23, Proposition 2.4.3.16.] this implies that the forgetful functor $\mathrm{Alg}_{{\mathbb{E}}_{\infty}}(\mathrm{Alg}_{\mathrm{Comm}}({\mathcal{C}}))\to\mathrm{Alg}_{\mathrm{Comm}}({\mathcal{C}})$ is an equivalence. So the symmetric monoidal forgetful functor $\mathrm{Alg}_{\mathrm{Comm}}({\mathcal{C}})\to{\mathcal{C}}$ lifts to a functor
@@
over ${\mathcal{C}}$ between monadic functors. Let ${\mathrm{S}}$ be the left adjoint of the forgetful functor $\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$. By [23, Corollary 4.7.3.16.] the functor $\rho$ is an equivalence if for every ${\mathrm{X}}\in{\mathcal{C}}$ the morphism $\psi:{\mathrm{S}}({\mathrm{X}})\to{\mathrm{S}}(\mathrm{Sym}({\mathrm{X}}))\to\mathrm{Sym}({\mathrm{X}})$ adjoint to the unit ${\mathrm{X}}\to\mathrm{Sym}({\mathrm{X}})$ is an equivalence. The counit ${\mathrm{S}}(\mathrm{Sym}({\mathrm{X}}))\simeq\coprod_{{\mathrm{n}}\geq 0}(\mathrm{Sym}({\mathrm{X}})^{\otimes{\mathrm{n}}})_{\Sigma{\mathrm{n}}}\to\mathrm{Sym}({\mathrm{X}})$ gives in each summand the multiplication of $\mathrm{Sym}({\mathrm{X}}).$ For every ${\mathrm{n}}\geq 0$ the map
$({\mathrm{X}}^{\otimes{\mathrm{n}}})_{\Sigma{\mathrm{n}}}\to(\mathrm{Sym}({\mathrm{X}})^{\otimes{\mathrm{n}}})_{\Sigma{\mathrm{n}}}\to\mathrm{Sym}({\mathrm{X}})\simeq\coprod_{{\mathrm{n}}\geq 0}({\mathrm{X}}^{\otimes{\mathrm{n}}})_{\Sigma{\mathrm{n}}},$
@@
${\mathrm{coAlg}}_{{\mathrm{coComm}}}({\mathcal{C}})\simeq{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}).$
@@ 2.6. Trivial algebras
${\mathrm{triv}}_{\mathcal{Q}}:{\mathcal{C}}\simeq{\mathrm{coAlg}}_{{\mathrm{triv}}}({\mathcal{C}})\to{\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}}).$
@@
$\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\xrightarrow{{\mathrm{Bar}}_{\mathcal{O}}(-)}\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathcal{C}})\xrightarrow{\mathrm{colim}}{\mathcal{C}}$
such that for any ${\mathrm{X}}\in\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})$ and $[{\mathrm{n}}]\in\Delta$ the object ${\mathrm{Bar}}_{\mathcal{O}}({\mathrm{X}})_{\mathrm{n}}$ is ${\mathcal{O}}\circ...\circ{\mathcal{O}}\circ{\mathrm{X}}$.
@@
Let ${\mathcal{C}}$ be a symmetric monoidal $\infty$-category such that ${\mathcal{C}}$ has small limits and the symmetric monoidal structure is compatible with the final object. Let ${\mathrm{triv}}\to{\mathcal{Q}}$ be a coaugmented $\infty$-cooperad in ${\mathcal{C}}$. Then ${\mathrm{triv}}_{\mathcal{Q}}:{\mathcal{C}}\to{\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}})$ has a right adjoint
${\mathrm{triv}}*^{\mathcal{Q}}(-):{\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}})\to{\mathcal{C}}$
@@
${\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}})\xrightarrow{{\mathrm{Cobar}}_{\mathcal{Q}}(-)}\mathrm{Fun}(\Delta,{\mathcal{C}})\xrightarrow{\lim}{\mathcal{C}}$
such that for any ${\mathrm{X}}\in{\mathrm{coAlg}}_{\mathcal{Q}}({\mathcal{C}})$ and $[{\mathrm{n}}]\in\Delta$ the object ${\mathrm{Cobar}}_{\mathcal{O}}({\mathrm{X}})_{\mathrm{n}}$ is ${\mathcal{O}}*...*{\mathcal{O}}*{\mathrm{X}}$.
@@
Let ${\mathcal{D}}$ be an $\infty$-category bitensored over monoidal $\infty$-categories ${\mathcal{C}},{\mathcal{E}}$, ${\mathrm{A}}$ an algebra in ${\mathcal{C}}$, ${\mathrm{B}}$ an algebra in ${\mathcal{E}}$, ${\mathrm{M}}$ an ${\mathrm{A}},{\mathrm{B}}$-bimodule, ${\mathrm{X}}$ a right ${\mathrm{A}}$-module and ${\mathrm{Y}}$ a left ${\mathrm{B}}$-module. By [23, Theorem 4.4.2.8.] there is a simplicial object ${\mathrm{Bar}}^{\mathcal{D}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},{\mathrm{M}},{\mathrm{Y}})$ in ${\mathcal{D}}$ such that
${\mathrm{Bar}}^{\mathcal{D}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},{\mathrm{M}},{\mathrm{Y}})_{\mathrm{n}}\simeq{\mathrm{X}}\otimes{\mathrm{A}}^{\otimes{\mathrm{n}}}\otimes{\mathrm{M}}\otimes{\mathrm{B}}^{\otimes{\mathrm{n}}}\otimes{\mathrm{Y}}.$
@@
By [23, Example 4.7.2.7.] the simplicial object ${\mathrm{Bar}}^{\mathcal{D}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{A}},{\mathrm{M}},{\mathrm{B}})$ in ${\mathcal{D}}$ canonically splits, and lifts and extends to an augmented simplicial object in ${{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})$ that exhibits ${\mathrm{M}}$ as ${\mathrm{A}}\otimes_{\mathrm{A}}{\mathrm{M}}\otimes_{\mathrm{B}}{\mathrm{A}}$ and sends $[{\mathrm{n}}]\in\Delta$ to the free ${\mathrm{A}},{\mathrm{B}}$-bimodule on ${\mathrm{A}}^{\otimes{\mathrm{n}}}\otimes{\mathrm{M}}\otimes{\mathrm{B}}^{\otimes{\mathrm{n}}}$.
Let ${\mathcal{D}}$ be an $\infty$-category lax bitensored over non-symmetric $\infty$-operads ${\mathcal{C}},{\mathcal{E}}$. By [19, Notation 3.108.] there is an enveloping $\infty$-category ${\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})$ bitensored over monoidal $\infty$-categories ${\mathrm{Env}}({\mathcal{C}}),{\mathrm{Env}}({\mathcal{E}})$, which comes equipped with embeddings ${\mathcal{C}}\subset{\mathrm{Env}}({\mathcal{C}}),{\mathcal{E}}\subset{\mathrm{Env}}({\mathcal{E}})$ of non-symmetric $\infty$-operads and an embedding ${\mathcal{D}}\subset{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})$. Every object of ${\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})$ is of the form ${\mathrm{V}}_{1}\otimes...\otimes{\mathrm{V}}_{\mathrm{n}}\otimes{\mathrm{X}}\otimes{\mathrm{W}}_{1}\otimes...\otimes{\mathrm{W}}_{\mathrm{k}}$ for ${\mathrm{V}}_{1},...,{\mathrm{V}}_{\mathrm{n}}\in{\mathcal{C}},{\mathrm{X}}\in{\mathcal{D}},{\mathrm{W}}_{1},...,{\mathrm{W}}_{\mathrm{k}}\in{\mathcal{E}}$ and ${\mathrm{k}},{\mathrm{n}}\geq 0.$ By [19, Lemma 3.103.] the embedding ${\mathcal{D}}\subset{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})$ admits a left adjoint ${\mathrm{L}}$ (not compatible with the biactions) that sends an object ${\mathrm{V}}_{1}\otimes...\otimes{\mathrm{V}}_{\mathrm{n}}\otimes{\mathrm{X}}\otimes{\mathrm{W}}_{1}\otimes...\otimes{\mathrm{W}}_{\mathrm{k}}$ for ${\mathrm{V}}_{1},...,{\mathrm{V}}_{\mathrm{n}}\in{\mathcal{C}},{\mathrm{X}}\in{\mathcal{D}},{\mathrm{W}}_{1},...,{\mathrm{W}}_{\mathrm{k}}\in{\mathcal{E}}$ to ${\mathrm{V}}_{1}\otimes...\otimes{\mathrm{V}}_{\mathrm{n}}\otimes{\mathrm{X}}\otimes{\mathrm{W}}_{1}\otimes...\otimes{\mathrm{W}}_{\mathrm{k}}$, where for the latter object the biaction of ${\mathcal{C}},{\mathcal{E}}$ on ${\mathcal{D}}$ is used.
@@
Let ${\mathcal{D}}$ be an $\infty$-category bitensored over monoidal $\infty$-categories ${\mathcal{C}},{\mathcal{E}}$, ${\mathrm{A}}$ an algebra in ${\mathcal{C}}$, ${\mathrm{B}}$ an algebra in ${\mathcal{E}}$, ${\mathrm{M}}$ an ${\mathrm{A}},{\mathrm{B}}$-bimodule, ${\mathrm{X}}$ a right ${\mathrm{A}}$-module and ${\mathrm{Y}}$ a left ${\mathrm{B}}$-module. Then ${\mathrm{Bar}}^{{\mathcal{D}}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})$ factors as
${{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})\subset{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))\xrightarrow{{\mathrm{Bar}}^{{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})}\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))$

### v2

@@ 2.5. Hopf $\infty$-operads
$\Gamma:\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})_{/{\mathbb{1}}}^{\prime}\to\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nun}}}({\mathcal{C}})$
that takes the fiber of the augmentation in $\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nun}}}({\mathcal{C}}).$ For any ${\mathrm{X}}\in\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nun}}}({\mathcal{C}})$ the unit
@@
where the vertical functors are fully faithful and the bottom functor is an equivalence. Since ${\mathcal{C}}\subset{\mathcal{D}}$ is closed under retracts, ${\mathrm{A}}\in\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nun}}}({\mathcal{D}})$ belongs to $\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nun}}}({\mathcal{C}})$ if its image $\bar{{\mathrm{F}}}({\mathrm{A}})\simeq{\mathrm{A}}\oplus{\mathbb{1}}$ belongs to $\mathrm{Alg}_{\mathrm{Comm}}({\mathcal{C}})_{/{\mathbb{1}}}$. So the top functor is an equivalence.
@@
    There is a canonical embedding $\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\hookrightarrow\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}}({\mathcal{C}})_{{\mathbb{1}}/}$ sending $\mathrm{X}$} to ${\mathrm{X}}\oplus{\mathbb{1}}.$
@@
    If $\mathcal{C}$} admits cofibers, the embedding $\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\hookrightarrow\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}}({\mathcal{C}})_{{\mathbb{1}}/}$ admits a left adjoint that takes the cofiber of the coaugmentation.
@@
    The embedding $\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\hookrightarrow\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}}({\mathcal{C}})_{{\mathbb{1}}/}$ is an equivalence if $\mathcal{C}$} is stable or idempotent complete and additive.
The next proposition is [23, Theorem 8.9.]:
@@
[…]
@@
${\mathrm{Sym}}:=\mathrm{Comm}\circ(-):{\mathcal{C}}\to\mathrm{Alg}_{\mathrm{Comm}}({\mathcal{C}}),\ {\mathrm{X}}\mapsto\mathrm{Comm}\circ{\mathrm{X}}\simeq\coprod_{{n}\geq 0}({\mathrm{X}}^{\otimes{n}})_{\Sigma{n}}.$
By Proposition 2.55 the forgetful functor $\mathrm{Alg}_{\mathrm{Comm}}({\mathcal{C}})\to{\mathcal{C}}$ refines to a symmetric monoidal functor since $\mathrm{Comm}$ is a Hopf $\infty$-operad in $\mathcal{C}$}. We start with proving that the symmetric monoidal structure on $\mathrm{Alg}_{\mathrm{Comm}}({\mathcal{C}})$ is cocartesian. The tensor unit $\mathbb{1}$} of $\mathrm{Alg}_{\mathrm{Comm}}({\mathcal{C}})$ is initial since the unique morphism $\alpha:{\mathrm{Sym}}(\emptyset)\to{\mathbb{1}}$ is an equivalence: indeed, the morphism $\alpha$ factors as ${\mathrm{Sym}}(\emptyset)\to{\mathrm{Sym}}({\mathbb{1}})\xrightarrow{\mu}{\mathbb{1}},$ where the counit $\mu$ is induced by the multiplication of $\mathbb{1}$}. So $\alpha$ is the 1-fold multiplication ${\mathbb{1}}^{\otimes 1}\to{\mathbb{1}}$ of the $\mathrm{Comm}$-algebra $\mathbb{1}$} and so the identity by unitality of the algebra.
@@
$\alpha:{\mathrm{Sym}}({\mathrm{X}}\coprod{\mathrm{Y}})\xrightarrow{{\mathrm{Sym}}(\beta)}{\mathrm{Sym}}({\mathrm{Sym}}({\mathrm{X}})\otimes{\mathrm{Sym}}({\mathrm{Y}}))\xrightarrow{\mu}{\mathrm{Sym}}({\mathrm{X}})\otimes{\mathrm{Sym}}({\mathrm{Y}}),$
@@
By [39, Proposition 2.4.3.16.] this implies that the forgetful functor $\mathrm{Alg}_{{\mathbb{E}}_{\infty}}(\mathrm{Alg}_{\mathrm{Comm}}({\mathcal{C}}))\to\mathrm{Alg}_{\mathrm{Comm}}({\mathcal{C}})$ is an equivalence. So the symmetric monoidal forgetful functor $\mathrm{Alg}_{\mathrm{Comm}}({\mathcal{C}})\to{\mathcal{C}}$ lifts to a functor
@@
over $\mathcal{C}$} between monadic functors. Let $\mathrm{S}$} be the left adjoint of the forgetful functor $\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$. By [39, Corollary 4.7.3.16.] the functor $\rho$ is an equivalence if for every ${\mathrm{X}}\in{\mathcal{C}}$ the morphism $\psi:{\mathrm{S}}({\mathrm{X}})\to{\mathrm{S}}({\mathrm{Sym}}({\mathrm{X}}))\to{\mathrm{Sym}}({\mathrm{X}})$ adjoint to the unit ${\mathrm{X}}\to{\mathrm{Sym}}({\mathrm{X}})$ is an equivalence. The counit ${\mathrm{S}}({\mathrm{Sym}}({\mathrm{X}}))\simeq\coprod_{{n}\geq 0}({\mathrm{Sym}}({\mathrm{X}})^{\otimes{n}})_{\Sigma{n}}\to{\mathrm{Sym}}({\mathrm{X}})$ gives in each summand the multiplication of ${\mathrm{Sym}}({\mathrm{X}}).$ For every ${n}\geq 0$ the map
$({\mathrm{X}}^{\otimes{n}})_{\Sigma{n}}\to({\mathrm{Sym}}({\mathrm{X}})^{\otimes{n}})_{\Sigma{n}}\to{\mathrm{Sym}}({\mathrm{X}})\simeq\coprod_{{n}\geq 0}({\mathrm{X}}^{\otimes{n}})_{\Sigma{n}},$
@@
$\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}}({\mathcal{C}})\simeq\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}).$
@@ 2.6. Trivial algebras
${\mathrm{triv}}_{\mathcal{Q}}:{\mathcal{C}}\simeq\mathrm{co}\mathrm{Alg}_{{\mathrm{triv}}}({\mathcal{C}})\to\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}}).$
@@
$\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})\xrightarrow{{\mathrm{Barc}}_{\mathcal{O}}(-)}\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathcal{C}})\xrightarrow{\mathrm{colim}}{\mathcal{C}}$
such that for any ${\mathrm{X}}\in\mathrm{Alg}_{\mathcal{O}}({\mathcal{C}})$ and ${n}]\in\Delta$ the object ${\mathrm{Barc}}_{\mathcal{O}}({\mathrm{X}})_{n}$ is ${\mathcal{O}}\circ...\circ{\mathcal{O}}\circ{\mathrm{X}}$.
@@
Let $\mathcal{C}$} be a symmetric monoidal $\infty$-category such that $\mathcal{C}$} has small limits and the symmetric monoidal structure is compatible with the final object. Let ${\mathrm{triv}}\to{\mathcal{Q}}$ be a coaugmented $\infty$-cooperad in $\mathcal{C}$}. Then ${\mathrm{triv}}_{\mathcal{Q}}:{\mathcal{C}}\to\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}})$ has a right adjoint
${\mathrm{triv}}*^{\mathcal{Q}}(-):\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}})\to{\mathcal{C}}$
@@
$\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}})\xrightarrow{{\mathrm{Cobar}}_{\mathcal{Q}}(-)}\mathrm{Fun}(\Delta,{\mathcal{C}})\xrightarrow{\lim}{\mathcal{C}}$
such that for any ${\mathrm{X}}\in\mathrm{co}\mathrm{Alg}_{\mathcal{Q}}({\mathcal{C}})$ and ${n}]\in\Delta$ the object ${\mathrm{Cobar}}_{\mathcal{O}}({\mathrm{X}})_{n}$ is ${\mathcal{O}}*...*{\mathcal{O}}*{\mathrm{X}}$.
@@
Let $\mathcal{D}$} be an $\infty$-category bitensored over monoidal $\infty$-categories $\mathcal{C}$$\mathcal{E}$}, $\mathrm{A}$} an algebra in $\mathcal{C}$}, $\mathrm{B}$} an algebra in $\mathcal{E}$}, $\mathrm{M}$} an ${\mathrm{A}},{\mathrm{B}}$-bimodule, $\mathrm{X}$} a right $\mathrm{A}$}-module and $\mathrm{Y}$} a left $\mathrm{B}$}-module. By [39, Theorem 4.4.2.8.] there is a simplicial object ${\mathrm{Barc}}^{\mathcal{D}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},{\mathrm{M}},{\mathrm{Y}})$ in $\mathcal{D}$} such that
${\mathrm{Barc}}^{\mathcal{D}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},{\mathrm{M}},{\mathrm{Y}})_{n}\simeq{\mathrm{X}}\otimes{\mathrm{A}}^{\otimes{n}}\otimes{\mathrm{M}}\otimes{\mathrm{B}}^{\otimes{n}}\otimes{\mathrm{Y}}.$
@@
By [39, Example 4.7.2.7.] the simplicial object ${\mathrm{Barc}}^{\mathcal{D}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{A}},{\mathrm{M}},{\mathrm{B}})$ in $\mathcal{D}$} canonically splits, and lifts and extends to an augmented simplicial object in ${{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})$ that exhibits $\mathrm{M}$} as ${\mathrm{A}}\otimes_{\mathrm{A}}{\mathrm{M}}\otimes_{\mathrm{B}}{\mathrm{A}}$ and sends ${n}]\in\Delta$ to the free ${\mathrm{A}},{\mathrm{B}}$-bimodule on ${\mathrm{A}}^{\otimes{n}}\otimes{\mathrm{M}}\otimes{\mathrm{B}}^{\otimes{n}}$.
Let $\mathcal{D}$} be an $\infty$-category lax bitensored over non-symmetric $\infty$-operads $\mathcal{C}$$\mathcal{E}$}. By [27, Definition 2.96.], [28, Proposition 2.95.] there is an enveloping $\infty$-category ${\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})$ bitensored over monoidal $\infty$-categories ${\mathrm{Env}}({\mathcal{C}}),{\mathrm{Env}}({\mathcal{E}})$, which comes equipped with embeddings ${\mathcal{C}}\subset{\mathrm{Env}}({\mathcal{C}}),{\mathcal{E}}\subset{\mathrm{Env}}({\mathcal{E}})$ of non-symmetric $\infty$-operads and an embedding ${\mathcal{D}}\subset{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})$. Every object of ${\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})$ is of the form ${\mathrm{V}}_{1}\otimes...\otimes{\mathrm{V}}_{n}\otimes{\mathrm{X}}\otimes{\mathrm{W}}_{1}\otimes...\otimes{\mathrm{W}}_{k}$ for ${\mathrm{V}}_{1},...,{\mathrm{V}}_{n}\in{\mathcal{C}},{\mathrm{X}}\in{\mathcal{D}},{\mathrm{W}}_{1},...,{\mathrm{W}}_{k}\in{\mathcal{E}}$ and ${k},{n}\geq 0.$ By [25, Lemma 3.103.] the embedding ${\mathcal{D}}\subset{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})$ admits a left adjoint $\mathrm{L}$} (not compatible with the biactions) that sends an object ${\mathrm{V}}_{1}\otimes...\otimes{\mathrm{V}}_{n}\otimes{\mathrm{X}}\otimes{\mathrm{W}}_{1}\otimes...\otimes{\mathrm{W}}_{k}$ for ${\mathrm{V}}_{1},...,{\mathrm{V}}_{n}\in{\mathcal{C}},{\mathrm{X}}\in{\mathcal{D}},{\mathrm{W}}_{1},...,{\mathrm{W}}_{k}\in{\mathcal{E}}$ to ${\mathrm{V}}_{1}\otimes...\otimes{\mathrm{V}}_{n}\otimes{\mathrm{X}}\otimes{\mathrm{W}}_{1}\otimes...\otimes{\mathrm{W}}_{k}$, where for the latter object the biaction of $\mathcal{C}$$\mathcal{E}$} on $\mathcal{D}$} is used.
@@
Let $\mathcal{D}$} be an $\infty$-category bitensored over monoidal $\infty$-categories $\mathcal{C}$$\mathcal{E}$}, $\mathrm{A}$} an algebra in $\mathcal{C}$}, $\mathrm{B}$} an algebra in $\mathcal{E}$}, $\mathrm{M}$} an ${\mathrm{A}},{\mathrm{B}}$-bimodule, $\mathrm{X}$} a right $\mathrm{A}$}-module and $\mathrm{Y}$} a left $\mathrm{B}$}-module. Then ${\mathrm{Barc}}^{{\mathcal{D}}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})$ factors as
${{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})\subset{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))\xrightarrow{{\mathrm{Barc}}^{{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})}\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))$